OpenStudy (anonymous):

prove if a<=0<=epsilon, for every epsilon>0, then a=0

OpenStudy (anonymous):

Hmm, are you sure it's not \[0 \le a \le \epsilon\], instead of \[a \le 0 \le \epsilon\]?